OpenStudy (anonymous):

Solve the equation 72+4-14d =36

OpenStudy (zzr0ck3r):

subtract 76 from both sides then divide by -14

OpenStudy (anonymous):

72 + 4 - 14d=36 36 - 4 = 32 Soooo: 72+14d=32 36 - 72= -40 Then you get: 14d= -40 -40 divided by 14d = -2.85

OpenStudy (luigi0210):

@brriiarr No doing others work or giving out answers

OpenStudy (zzr0ck3r):

he showed plenty of steps, I see the mods doing this all the time. So I am very confused on what is ok and what is not....

OpenStudy (anonymous):

Sorry. I just assumed since like most of the people that answer questions on here give out the answers, it was an okay thing. Especially since, you know, they all get away with it.

OpenStudy (anonymous):

I wasn't doing this just so i could get the answer, i was really confused on it so i needed to know how it was done.

OpenStudy (anonymous):

Don't worry about it, you're fine :)

OpenStudy (anonymous):

Ok thanks :)